5. The first term of an AP is 5, the last term is 45 and the sum is 400. Find the number of terms and the common difference.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given that, first term, a = 5 last term, l = 45 Sum of the AP, Sn = 400 As we know, the sum of AP formula is; Sn = n/2 (a+l) 400 = n/2(5+45) 400 = n/2(50) Number of terms, n =16 As we know, the last term of AP series can be written as; l = a+(n −1)d 45 = 5 +(16 −1)d 40 = 15d Common difference, d = 40/15 = 8/3